40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Message ID: 1311     Entry time: Fri Feb 18 20:07:44 2011
Author: Zach 
Type: Laser 
Category: GYRO 
Subject: MZ measurement in progress 

 [Alastair, Zach]

We finished cleaning the in-vac stuff, installed the mirrors in the MZ configuration, injected a beam, aligned it for best contrast, put the output on a PD, and then pumped down. We disconnected the pump once we got down to the level that Alastair had achieved previously (~a few x 10-4 torr) and the pressure rose to a few hundredths of a torr.

The output of the interferometer throughout the process was a slow sine wave, which we figured was due to the small changes in cavity length from the pumping/leaking.

The channel is being acquired. More analysis and a better explanation of the setup to come.

ELOG V3.1.3-